# Federated Learning Privacy: Attacks, Defenses, Applications, and Policy Landscape — A Survey

### v1

@@ 4.3. Application of FL in IoT or at the Edge
The Internet of Things (IoT) is the current concept of ubiquitous internet connectivity of all devices (refrigerators, toasters, doorbells, street cameras, …, the list is unending). As these devices suggest the availability of data from these devices from anywhere, the privacy and security of these devices and the data is paramount. However, there is strong interest from companies and individuals to exploit the available data for constructive applications to benefit more than just the immediate use-case and users. One prominent example is the Google keyboard (Gboard) prediction (Hard et al., 2018), where Google improves the performance of their mobile keyboard prediction by gaining information from user interaction. However, this example highlights information privacy in that improvements to this technology will benefit all users but users are willing to share only if their privacy is assured. Therefore, FL techniques that exploit data and maintain user and data privacy are attractive for IoT applications (Zhang et al., 2021).
@@
In such scenarios and related application domains, it is well understood that communication and computing resources are two of the bottlenecks for federated learning in real-life applications. The devices are often used for tasks other than FL, so the FL task usually competes with these other ones for computing resources, making scheduling and coordinating even more complicated. Asynchronous strategies like PAPAYA(Huba et al., 2022) appear to address such issues. Other scenarios include heterogeneity in the computing resources (Diao et al., 2021) or drift in local device clocks (Koo et al., 2009), leading to asynchrony caused by varying compute and resulting in delays in the aggregation step. Methods to account for or dynamically adapt to changing available resources will greatly enhance the utility of FL techniques in these scenarios. The increased need for bandwidth for FL iterations has led to significant research efforts in reducing it. Methods combining optimizers such as FedAvg with sparsification and/or quantization of model updates to a small number of bits have demonstrated significant reductions in communication overhead. However, it still remains an open research question how communication overheads can be further reduced, and whether any of these methods or their combinations can come close to providing optimal trade-offs between communication and accuracy in federated learning. For example, the Fed-ET approach is promising where smaller models are communicated (Cho et al., 2022). Another class of approach to maximizing bandwidth usage in FL approaches include device selection (Perazzone et al., 2022). While efforts in these situations are motivated by increasing communications efficiency, device selection methods can potentially provide device privacy in terms of revealing which nodes are contributing to FL process. Moreover, there has also been recent work that allows for the control information exchange to limit communications, but can similarly be used to provide some privacy to node utility to FL algorithms (Wang et al., 2023).
@@
Additionally, IoT has been promoted as a use-case for edge networks where networks are bandwidth-limited and storage and compute is also limited. These systems operate in a variety of environments often characterized by having stringent operational requirements despite having to cope with sophisticated adversarial influences and constraints in both computational and network resources. These environments include resource-constraints inhibiting information exchange to other nodes, specifically a central node with which centralized computation and decisions can be conducted. Federated approaches have been proposed for use in sensor networks, swarms of unmanned aerial vehicle and on coalition networks for shared situational awareness. They have also been proposed for FL at the edge by reducing the network overhead incurred by such learning processes (Wang et al., 2019b). The promise of IoT and edge computing will continue to be revealed, with FL techniques potentially as a main driver of the technologies.
@@ 5. Privacy-Related Policies
Most of the existing regulations on safeguarding consumers privacy, both nationally and internationally were conceived in a pre-AI/ML era and have not been able to keep up with the rapid evolution of AI/Ml and their privacy implications. The European Union’s General Data Protection Regulation (”GDPR”) (GDP, [n. d.]) governs data protection and privacy for all citizens within the EU and imposes strict obligations on data controllers and processors. GDPR contains the principle of data minimization in article 5.1 (c) which requires that the data that are not necessary to achieve the intended purpose cannot be lawfully collected, stored or otherwise processed. Since federated learning restricts the transfer of raw data to a central location, it provides more compliance with the principle of data minimization than other approaches that do require the transfer of all raw date to a central location to train a model. Federated learning also protects the collected data from unwanted processing that are not compliant with the collection purpose satisfying the principle of purpose limitation, as described in article 5.1.(b).
@@
In the United States, the state of Privacy and AI laws is tracked by the EPIC project which shows recent large number of states effective privacy laws (EPI, [n. d.]). Each of these privacy laws has specific requirements when accessing and processing personal data. Even FTC is enforcing the same privacy rights for companies using AI (FTC, [n. d.]). At the federal level, sector-specific privacy polices in the US such as Graham-Leach-Bliley Act (GLBA) (GLB, [n. d.]) for financial data or Health Insurance Portability and Accountability Act (HIPAA) (HIP, [n. d.]) for medical date provide consumers with limited protection scope when it comes to using AI and ML. Recently, US introduced H.R. 8152-The American Data Privacy and Protection Act (HR8, [n. d.])-to catch up with various amendments proposed in EU on AI/ML workloads by extending the covered algorithm to AI/ML space. Some states also patched their existing policies such as California Privacy Rights Act (CPRA) (CPR, [n. d.]) amendments to the CCPA (CCP, [n. d.]) or Texas Data Privacy and Security Act (TDPSA) (tex, [n. d.]).
@@
GenAI’s remarkable capacity to analyze data and make complex analyses further amplifies privacy concerns. There is clear a privacy gaps in existing privacy policies as it relates to AI systems. Most of these policies and regulations were proposed pre-AI and even newer ones per Gen AI where there has been less information on privacy implications of these systems. Such gap may pose significant challenges to privacy laws and they will need further amendments to address emerging privacy concerns from Gen AI.
@@
The recent AI Foundation Model Transparency Act (AIF, [n. d.]) mandates FTC and NIST to establish standards for data sharing by foundation model deployers. This legislation goal is to empower consumers to make well informed decisions when they interact with AI. The Algorithmic Accountability Act of 2023 (AIA, [n. d.]) also requires companies to assess the impacts of the AI systems on consumer privacy and give users choices when they interact with AI systems.
@@
These regulations mostly require keeping the data in its location and restrict its processing based on the intended collection purposes. Federated learning is a potential solution to facilitate compliance with these privacy laws.
@@ 6.1. Data Reconstruction Attacks
| Attack class | Summary | Sub-category | Citation | Requirements | |||||||||||||
| Data reconstruction | Attacker aims to directly reconstruct the client private data or a representation of the private data. | Gradient inversion | (Zhu et al., 2019; Zhao et al., 2020; Geiping et al., 2020; Yin et al., 2021a) |
(Zhao et al., 2023b; Fowl et al., 2022; Boenisch et al., 2023) |
| GAN-based | (Hitaj et al., 2017; Wang et al., 2019a) |
| Other | (Wen et al., 2022; Pasquini et al., 2022; Lam et al., 2021) |
| Membership inference |
(Song et al., 2019; Li et al., 2022a; Nasr et al., 2019; Zari et al., 2021; Melis et al., 2019) |
| Property inference |
(Fredrikson et al., 2015; Mehnaz et al., 2022; Dibbo et al., 2023; Ganju et al., 2018) |
| Model extraction |
(Tramèr et al., 2016; Wang and Gong, 2018; Orekondy et al., 2019) |
@@
Optimization (a.k.a. gradient inversion) approaches have shown great success in leaking data from individual updates, especially with smaller batch sizes. These attacks typically operate under the threat model of an honest-but-curious server or an external attacker that has access to the model and individual gradients from each client. With only this information, the attacker initializes some dummy data and computes the gradient of that data on the model.
An optimizer minimizes the difference between the generated gradient and the ground truth gradient (benign client update). Here is the dummy data, is the reconstructed data, is the loss function, is the label, and is the model parameters.
More recent optimization approaches (Geiping et al., 2020; Yin et al., 2021a) work under the assumption that user labels are known prior to optimization. Typically, these labels are directly retrieved through a zero-shot solution without using optimization approaches (Zhao et al., 2020; Yin et al., 2021a). Furthermore, regularizers and strong image priors specific to image data are often used to guide optimization results (Geiping et al., 2020; Yin et al., 2021a; Hatamizadeh et al., 2023; Usynin et al., 2023). These can also result in image artifacts typical of an image class, but not in the actual training image. These approaches have shown surprising success with image data on smaller batch sizes. However, as batch sizes increase, the fraction of images recovered decreases along with the reconstruction quality and the number of iterations required for the optimization also increases. One reason stated by (Zhu et al., 2019) was that regardless of the order of images in the batch, the gradient will remain the same. Having multiple possible permutations then makes the optimization more difficult. Another fundamental reason is that a larger batch size means more images and more variables for optimization.
Both secure aggregation and FedAvg also pose a particularly difficult challenge for optimization attacks. For secure aggregation, the aggregated updated can be thought of as large-batch consisting of the batch images from all clients. This directly exacerbates the difficulty of reconstructing larger batch sizes. FedAvg, on the other hand, adds another layer of unknowns for the optimization since the intermediate model updates are unknown to the attacker. While (Geiping et al., 2020) discuss attacking FedAvg, the attack is only shown with a very small local dataset size of up to 8 images on CIFAR-10. Furthermore, it needs to be investigated how different model architectures, such as transformers (Hatamizadeh et al., 2022), impact the success of gradient inversion attacks.
@@
Linear layer leakage attacks are a sub-class of analytic attacks that modify FC layers to leak inputs. Using the weight and bias gradients of an FC layer to leak inputs was discussed in (Phong et al., 2017; Fan et al., 2020). When only a single image activates a neuron in a fully connected layer, the input to that layer can be directly computed using the resulting gradients as
where is the activated neuron, is the input that activates neuron , and , are the weight gradient and bias gradient of the neuron respectively. This idea forms the basis for several reconstruction attacks (Boenisch et al., 2023; Fowl et al., 2022). Figure 4 shows the basic process of leaking images through an FC layer.
When the fully-connected layer is placed at the start of a network, the data reconstructed from the the layer would be the input data. This reconstruction is exact, as opposed to the optimization approaches which function as estimations. However, inputs are only reconstructed exactly when a single data sample activates that neuron. If more than one input activates the neuron, the weight and bias gradients of these inputs will contribute to the batch gradient. When the gradient division of Equation 6 is done to retrieve the input, the resulting reconstruction would be a combination of all contributing images, a case of failed attack.
To alleviate this problem, (Fowl et al., 2022; Boenisch et al., 2023) use malicious modification of the parameters in the FC layer. For (Boenisch et al., 2023), trap weights were introduced, initializing the weights randomly to be half positive, half negative. In order to ensure that neuron activation is less common, the negative weights come from a larger negative magnitude range than the positive weights. They also discuss the use of convolutional layers to push the input image forward, allowing the attack to function on models starting with convolutional layers followed by fully-connected layers. However, one of the main problems of the method lies with scalability. Even if the size of the FC layer increases proportionately with an increasing total number of images, the leakage rate decreases. On the other hand, Robbing the Fed (RtF) (Fowl et al., 2022) introduced another approach with higher leakage rate called “binning”, where the weights of the FC layer would measure some known continuous CDF of the input data such as image brightness. The bias for each neuron then serves as a different cutoff, allowing only inputs with a high enough value to activate it. The goal of this method would be that only one input activates each “bin”, where the bin is defined as the activated neuron with the largest cutoff (for ReLU, the largest negative bias)111The bin biases are set as negative. The weights are positive and so the negative bins are used to prevent ReLU activation.. For any case where only one input activates a bin, it can then be reconstructed as
@@
FedAvg. While the previous method works for FedSGD, for FedAvg, the model changes during local iterations and this prevents the reconstruction attack. As a result, (Fowl et al., 2022) proposed the sparse variant of the attack which uses an activation function with a double-sided threshold (e.g., Hardtanh) such as:
@@
Using this activation function, neuron activation will be sparse (i.e., images will only activate a single neuron). However, this range between 0 and 1 for the non-zero gradient is fixed for all neurons. Since RtF’s approach sets up neuron biases following a distribution of the images, the weights and biases of the FC layer will need to be adjusted to follow the new non-zero gradient range. This requires scaling the magnitude of these parameters based on the distance between the subsequent neuron biases. Consider that the weights originally measure average pixel brightness. In this case, all the weights would originally be set to , where is the total number of pixels. Then, the weights and biases are rescaled as:
where and are the scaled weights and biases of neuron respectively and is the distance between adjacent biases in the original distribution. This process uses the same distribution as the FedSGD case to setup the initial biases, while incorporating the fixed range of the new activation function by scaling the parameters. In FedAvg, after the clients send the updated model parameters, the server computes a “gradient” as:
@@
This FedAvg attack (Fowl et al., 2022) works well in the context of single client attacks. However, attacking secure aggregated gradients becomes a problem as the total images in the gradient increases multiplicative to the number of clients. Particularly, (Zhao et al., 2023b) finds that numerical precision poses a problem for the FedAvg linear layer leakage attack when a large number of clients is attacked. Through the use of convolutional layers and a convolutional scaling factor (CSF), (Zhao et al., 2023b) is able to scale to an arbitrary number of clients while maintaining a high leakage rate. Furthermore, the CSF improves on the multiple-image reconstruction problem, only allowing images within the same local mini-batch to activate the same units. Notably, this allows the leakage rate on FedAvg attacks to be higher than in FedSGD (where FedAvg is typically harder to attack than FedSGD).
Linear layer leakage reconstruction quality is typically higher quality than optimization-based attacks and does not rely on the number of clients. However, it also requires malicious modification of the model parameters and/or architecture, resulting in more detectable attacks. Furthermore, the attack also adds a resource overhead to the model size. This problem is exacerbated by the number of clients as the attack layer sizes need to scale to keep the attack effective. This problem and the use of sparse tensors to alleviate the problem are discussed by (Zhao et al., 2023a).
@@
Generative networks were introduced to attack collaborative learning in (Hitaj et al., 2017; Wang et al., 2019a). However, contrary to optimization-based attacks and linear layer leakage, the attack does not aim to directly reconstruct exact training images, but instead reconstructs representations of a targeted class which the attacker does not hold any images of. In the attack, an adversary (a client in collaborative or federated learning training) trains a generative network in conjunction with the benign collaborative model. For a single training round, the attack process goes as follows:
@@
In (Wen et al., 2022), attackers attack a target image in an arbitrarily large batch by manipulating the model parameters to magnify the update of only the single image. Magnification is done in two steps: the magnification of the target class followed by a specific feature for a target image. In order to estimate a boundary for the features, the server uses several rounds of additional FL training for setup. This requires for the batch data to be the same across training rounds. In (Kariyappa et al., 2023), the attacker formulates the problem of data reconstruction as a blind source separation problem. Given the observation that inputs to a linear layer can be directly reconstructed through the gradients, solving for the inputs is similar to solving for a unknown variables given a set of linear equations.
These attacks are all able to break some levels of aggregation. (Pasquini et al., 2022; Lam et al., 2021) break aggregation itself to allow for other data reconstruction attacks to work with an aggregate setting. (Wen et al., 2022; Kariyappa et al., 2023) both function on both individual and aggregate updates. Each of these methods have their own strengths and weaknesses. (Pasquini et al., 2022) requires sending different models to different clients but allows an attacker to single out a single update. (Lam et al., 2021) requires multiple training rounds and additional knowledge on top of the client update. However, it can essentially completely disaggregate aggregate updates over time. (Kariyappa et al., 2023) can attack batch sizes up to 1024, but also requires malicious modification of the network parameters. (Wen et al., 2022) targets a single image in an arbitary update size, but requires malicious modification of network parameters along with a few rounds of setup.

### v2

@@ 4.3. Application of FL in IoT or at the Edge
The Internet of Things (IoT) is the current concept of ubiquitous internet connectivity of all devices (refrigerators, toasters, doorbells, street cameras, …, the list is unending). As the data from these devices scan come from anywhere, the privacy and security of these devices and the data is paramount. However, there is strong interest from companies and individuals to exploit the available data for constructive applications to benefit more than just the immediate use case and users. One prominent example is the Google keyboard (Gboard) prediction function (Hard et al., 2018), where Google improves the performance of their mobile keyboard prediction by gaining information from user interaction. This example highlights information privacy in that improvements to this technology will benefit all users, although users are only willing to contirbute if their privacy is ensured. Therefore, FL techniques that exploit data and maintain user and data privacy are attractive for IoT applications (Zhang et al., 2021b).
@@
In such scenarios and related domains, it is well understood that communication and computing resources are two of the bottlenecks in real-life FL applications. These devices are often used for tasks other than FL, so the FL task usually competes with these other ones for computing resources, making scheduling and coordinating even more complicated. Asynchronous strategies like PAPAYA (Huba et al., 2022) appear to address such issues. Other scenarios include heterogeneity in computing resources (Diao et al., 2021) or drift in local device clocks (Koo et al., 2009), leading to asynchrony caused by varying compute and resulting in delays in the aggregation step. Methods to account for or dynamically adapt to changing available resources will greatly enhance the utility of FL techniques in these scenarios. The increased need for bandwidth for FL iterations has led to significant research efforts in reducing it. Methods combining optimizers such as FedAvg with sparsification and/or quantization of model updates to a small number of bits have demonstrated significant reductions in communication overhead. However, it still remains an open research question on how communication overheads can be further reduced and whether any of these methods or their combinations can come close to providing optimal trade-offs between communication and accuracy. For example, the Fed-ET approach communicates using smaller models (Cho et al., 2022). Another class of approach to maximizing bandwidth usage in FL approaches include device selection (Perazzone et al., 2022). Although efforts in these situations are motivated by increasing communications efficiency, device selection methods can potentially provide device privacy in terms of revealing which nodes are contributing to the FL process. Moreover, there has also been recent work that allows control of the information exchange to limit communications, but can similarly be used to provide some privacy of node utility to FL algorithms (Wang et al., 2023).
@@
Additionally, IoT has been promoted as a use-case for edge networks where bandwidth, storage, and compute are limited. These systems operate in a variety of environments often characterized a stringent operational requirements, despite having to cope with sophisticated adversarial influences and constraints in both computational and network resources. These environments include resource-constraints inhibiting information exchange to other nodes, specifically a central node with which centralized computation and decisions can be conducted. Federated approaches have been proposed for use in sensor networks, swarms of unmanned aerial vehicles and on coalition networks for shared situational awareness. They have also been proposed for FL at the edge by reducing the network overhead incurred by such learning processes (Wang et al., 2019b). The promise of IoT and edge computing will continue to be revealed, with FL techniques potentially being a main driver of the technologies.
@@ 5. Privacy-Related Policies
Most of the existing regulations on safeguarding consumer privacy, both nationally and internationally, were conceived in a pre-AI/ML era. Currently, they have not been able to keep up with the rapid evolution of AI/Ml and their privacy implications. The General Data Protection Regulation (”GDPR”) (GDP, [n. d.]) from the European Union governs data protection and privacy for all citizens within the EU and imposes strict obligations on data controllers and processors. GDPR contains the principle of data minimization in article 5.1 (c), which requires that data which is not necessary to achieve the intended purpose cannot be collected, stored, or otherwise processed lawfully. Since FL restricts the transfer of raw data to a central location, it provides more compliance with the principle of data minimization than other approaches that require the transfer of all raw data to a central location for model training. FL also protects the collected data from unwanted processing that is not compliant with the collection purpose, satisfying the principle of purpose limitation, as described in article 5.1.(b).
@@
In the United States, the state of Privacy and AI laws is tracked by the EPIC project, which shows a recent large number of states that have effective privacy laws (EPI, [n. d.]). Each of these privacy laws has specific requirements for accessing and processing personal data. Even FTC is enforcing the same privacy rights for companies using AI (FTC, [n. d.]). At the federal level, sector-specific privacy polices in the US, such as Graham-Leach-Bliley Act (GLBA) (GLB, [n. d.]) for financial data or Health Insurance Portability and Accountability Act (HIPAA) (HIP, [n. d.]) for medical date provide consumers with limited protection scope when it comes to using AI and ML. Recently, the US introduced H.R. 8152-The American Data Privacy and Protection Act (HR8, [n. d.])-to catch up with various amendments proposed in EU on AI/ML workloads by extending the covered algorithm to AI/ML space. Some states also patched their existing policies such as California Privacy Rights Act (CPRA) (CPR, [n. d.]) amendments to the CCPA (CCP, [n. d.]) or the Texas Data Privacy and Security Act (TDPSA) (tex, [n. d.]).
@@
GenAI’s remarkable capability to analyze data and perform complex analysis further amplifies privacy concerns. There are clear privacy gaps in existing privacy policies as they relate to AI systems. Most of these policies and regulations were proposed pre-AI and even newer ones pre-Gen AI where there has been less information on privacy implications of these systems. Such gaps may pose significant challenges to privacy laws and they will need further amendments to address newer emerging privacy concerns.
@@
The recent AI Foundation Model Transparency Act (AIF, [n. d.]) mandates FTC and NIST to establish standards for data sharing by foundation model deployers. This legislation aims to empower consumers to make welliinformed decisions when they interact with AI. The Algorithmic Accountability Act of 2023 (AIA, [n. d.]) also requires companies to assess the impacts of the AI systems on consumer privacy and give users choices when they interact with AI systems.
@@
Future policy. The previous regulations mostly require keeping the data in its location and restrict its processing based on intended collection purposes. GL is a potential solution to facilitate compliance with these privacy laws. In the near future, we anticipate a growing alignment between technological advancements – especially in artificial intelligence (AI) and FL – and evolving regulatory frameworks. This alignment will likely follow several key pathways.
First, there is a need for regulation specialized for AI use cases and application domains. Rather than relying on adapting existing frameworks like GDPR or HIPAA, regulatory bodies are expected to introduce laws specifically tailored to AI. These regulations will address the entire AI model lifecycle, covering aspects such as data collection, model training, decision-making, and retraining. The goal will be to ensure greater accountability, fairness, and transparency in AI systems. Additionally, the rise of generative AI requires a stronger focus on regulating not just the input data but also the output of AI systems. Transparency around how these generative systems, including large language models, are trained, the datasets they rely on, and how their outputs might impact privacy or perpetuate or amplify biases will become increasingly important. Moreover, governments can mandate real-time monitoring and reporting tools to ensure that AI systems maintain compliance with privacy regulations as they evolve dynamically.
Second, FL is poised to emerge as a valuable tool for compliance. As its use becomes more widespread, regulators may view it as a method that aligns with privacy-by-design principles. FL could be explicitly incorporated into compliance standards, particularly those influenced by GDPR-like data minimization principles, i.e., limit the collection of personal information to what is essential for a defined objective (GDP, [n. d.]). By leveraging a decentralized approach, companies may be able to innovate while ensuring compliance with data privacy laws.
Overall, FL aligns well with privacy regulations like GDPR and HIPAA. It adheres to GDPR’s principles of data minimization and purpose limitation. By keeping sensitive data on local servers, it significantly reduces the amount of sensitive information being processed centrally. Only model updates, which are typically less sensitive than raw data, are shared with the central server. FL systems can also be designed to collect only the data necessary for the specific purpose of model training. Furthermore, individuals’ rights over their data is supported by allowing users to opt-in or opt-out of the training process at any time. This aligns well with GDPR’s emphasis on user control and consent. The decentralized nature of FL enhances data security, reducing the risk of large-scale data breaches. It embodies the principle of privacy by design which is a key requirement under GDPR.
Despite the above advantages, FL still faces some regulatory challenges. Data privacy laws differ across countries and regions, which can complicate global implementation of FL systems. Further, ensuring and demonstrating compliance across a distributed system can be complex and requires robust observability and documentation practices. The process of combining multiple models during aggregation can pose security risks, such as susceptibility to model poisoning attacks. In the next two sections, we discuss new and emerging privacy threats along with other defenses such as DP-SGD or secure aggregation which may be required to continue to protect user privacy in FL.
@@ 6.1. Data Reconstruction Attacks
Attack class Summary Sub-category Citation Requirements Data reconstruction Attacker aims to directly reconstruct the client private data or a representation of the private data. Typically, this involves analyzing gradient updates or partial model parameters that inadvertently reveal features of the original input. With sufficient information, the attacker can refine these reconstructions to near-original fidelity, posing severe privacy risks—particularly in sensitive domains like healthcare. Factors such as model complexity, dataset size, and update granularity can further exacerbate this threat. Gradient inversion (Zhu et al., 2019; Zhao et al., 2020; Geiping et al., 2020; Yin et al., 2021a) Access to model and individual updates. Linear layer leakage (Zhao et al., 2023b; Fowl et al., 2022; Boenisch et al., 2023b) Modification to the network parameters and/or architecture. GAN-based (Hitaj et al., 2017; Wang et al., 2019a) Clients target a specific class by sending malicious updates. Other (Wen et al., 2022; Pasquini et al., 2022; Lam et al., 2021) Varies based on attack. Attacker may modify the model parameters, know training info, or send different updates. Membership inference Attacker infers if a sample was used in the training of the model. (Song et al., 2019; Nasr et al., 2019; Zari et al., 2021; Melis et al., 2019) Most require access to training data from the overall distribution. Some attacks require white-box access to the model or the ability to send malicious updates. Property inference Attacker learns about sensitive properties within the training set (e.g., race, gender, age). (Fredrikson et al., 2015; Mehnaz et al., 2022; Dibbo et al., 2023; Ganju et al., 2018) Attacker has black-box access to the model and can access the output. They may also possess other attributes beyond the targeted one. Model extraction Attacker wants to steal functionality of a model. This can be the parameters, hyperparameters, etc. (Tramèr et al., 2016; Wang and Gong, 2018; Orekondy et al., 2019) Attacker needs black-box access to the model in order to query it and observe the outputs.
@@
Optimization (a.k.a. gradient inversion) approaches have shown great success in leaking data from individual updates, especially with smaller batch sizes. These attacks typically operate under the threat model of an honest-but-curious server or an external attacker that has access to the model and individual gradients from each client. With only this information, the attacker initializes some dummy data (i.e., with gaussian noise) and computes the gradient of the data.
An optimizer minimizes the difference between the generated gradient and the ground truth gradient (benign client update). In the above equation, is the dummy data, is the reconstructed data, is the loss function, is the label (which can be known beforehand or solved for (Zhao et al., 2020)), and is the model parameters.
More recent optimization approaches (Geiping et al., 2020; Yin et al., 2021a) work under the assumption that user labels are known prior to optimization. Typically, these labels are directly retrieved through a zero-shot solution without using optimization approaches (Zhao et al., 2020; Yin et al., 2021a). Furthermore, regularizers and strong image priors specific to image data are often used to guide optimization results (Geiping et al., 2020; Yin et al., 2021a; Hatamizadeh et al., 2023; Usynin et al., 2023). These can also result in image artifacts typical of an image class, but not in the actual training image. These approaches have shown surprising success with image data on smaller batch sizes. However, as batch sizes increase, the fraction of images recovered decreases along with the reconstruction quality and the number of iterations required for the optimization also increases. One reason stated by (Zhu et al., 2019) was that regardless of the order of images in the batch, the gradient will remain the same. Having multiple possible permutations then makes the optimization more difficult. Another fundamental reason is that a larger batch size means more images and more variables for optimization. Figure 1 shows Inverting Gradients (Geiping et al., 2020) reconstructions of batch size 8 and 16 on CIFAR-100, demonstrating this property of lowering quality with larger batch sizes.
Both secure aggregation and FedAvg also pose a particularly difficult challenge for optimization attacks. For secure aggregation, the aggregated updated can be thought of as large-batch consisting of the batch images from all clients. This directly exacerbates the difficulty of reconstructing larger batch sizes. FedAvg, on the other hand, adds another layer of unknowns for optimization since the intermediate model updates are unknown to the attacker. While (Geiping et al., 2020) discuss attacking FedAvg, the attack is only shown with a very small local dataset size of up to 8 images on CIFAR-10. Furthermore, it needs to be investigated how different model architectures, such as transformers (Hatamizadeh et al., 2022), impact the success of gradient inversion attacks.
@@
Linear layer leakage attacks are a sub-class of analytic attacks that modify FC layers to leak inputs. Using the weight and bias gradients of an FC layer to leak inputs was discussed in (Phong et al., 2017; Fan et al., 2020). When only a single image activates a neuron in a fully connected layer, the input to that layer can be directly reconstructed with:
where is the activated neuron, is the input that activates neuron , and , are the weight gradient and bias gradient of the neuron respectively. This idea forms the basis for several reconstruction attacks (Boenisch et al., 2023b; Fowl et al., 2022). Figure 3 shows the basic process of leaking images through an FC layer.
When the fully-connected layer is placed at the start of a network, the data reconstructed from that layer would be the input data. This reconstruction is exact, as opposed to the optimization approaches which function as estimations. However, inputs are only reconstructed exactly when a single data sample activates that neuron. If more than one input activates the neuron, the weight and bias gradients of these inputs will contribute to the batch gradient. When the gradient division of Equation 6 is done to retrieve the input, the resulting reconstruction would be a combination of all contributing images, a case of failed attack.
To alleviate this problem, (Fowl et al., 2022; Boenisch et al., 2023b; Zhao et al., 2023b) use malicious modification of the parameters in the FC layer. For (Boenisch et al., 2023b), trap weights were introduced, initializing the weights randomly to be half positive, half negative. In order to ensure that neuron activation is less common, the negative weights come from a larger negative magnitude range than the positive weights. They also discuss the use of convolutional layers to push the input image forward, allowing the attack to function on models starting with convolutional layers followed by fully-connected layers. However, one of the main problems of the method lies with scalability. Even if the size of the FC layer increases proportionately with an increasing total number of images, the leakage rate decreases. On the other hand, Robbing the Fed (RtF) (Fowl et al., 2022) introduced another approach with higher leakage rate called “binning”, where the weights of the FC layer would measure some known continuous CDF of the input data such as image brightness. The bias for each neuron then serves as a different cutoff, allowing only inputs with a high enough value to activate it. The goal of this method would be that only one input activates each “bin”, where the bin is defined as the activated neuron with the largest cutoff (for ReLU, the largest negative bias)222The bin biases are set as negative. The weights are positive and so the negative bins are used to prevent ReLU activation.. For any case where only one input activates a bin, it can then be reconstructed as
@@
FedAvg. While the previous method works for FedSGD, for FedAvg, the model changes during local iterations and this prevents the reconstruction attack. As a result, Fowl et al. (2022) proposed the sparse variant of the attack which uses an activation function with a double-sided threshold (e.g., Hardtanh) such as:
@@
Using this activation function, neuron activation will be sparse (i.e., images will only activate a single neuron). However, this range between 0 and 1 for the non-zero gradient is fixed for all neurons. Since RtF’s approach sets up neuron biases following a distribution of the images, the weights and biases of the FC layer will need to be adjusted to follow the new non-zero gradient range. This requires scaling the magnitude of these parameters based on the distance between the subsequent neuron biases. Consider that the weights originally measure the average pixel brightness. In this case, all weights would originally be set to , where is the total number of pixels. Then, the weights and biases are rescaled as:
where and are the scaled weights and biases of neuron , respectively, and is the distance between adjacent biases in the original distribution. This process uses the same distribution as the FedSGD case to setup the initial biases, while incorporating the fixed range of the new activation function by scaling the parameters. In FedAvg, after the clients send the updated model parameters, the server computes a “gradient” as:
@@
This FedAvg attack (Fowl et al., 2022) works well in the context of single client attacks. However, attacking secure aggregated gradients becomes a problem as the total number of images in the gradient increases multiplicatively to the number of clients. In a recent attack, Zhao et al. (2023b) (LOKI) finds that numerical precision poses a problem for the FedAvg linear layer leakage attack when a large number of clients are attacked. Through the use of convolutional layers and a convolutional scaling factor (CSF), LOKI is able to scale to an arbitrary number of clients while maintaining a high leakage rate. Furthermore, the CSF improves on the multiple-image reconstruction problem, only allowing images within the same local mini-batch to activate the same units. Notably, this allows the leakage rate of FedAvg attacks to surpass the leakage rate in FedSGD (where FedAvg is typically thought to be harder to attack than FedSGD). Figure 2 shows the reconstructions from LOKI on a total of 64 images from CIFAR-100 in FedAVG (split over 8 local iterations). LOKI achieves a high leakage rate, with images reconstructed with high quality.
Linear layer leakage reconstruction quality is typically higher quality than optimization-based attacks and is not impacted by the number of clients. However, it also requires malicious modification of the model parameters and/or architecture, resulting in more detectable attacks. Furthermore, the attack also adds a resource overhead to the total model size. This problem is exacerbated by secure aggregation, as the size of the attack layer needs to scale to keep the attack effective. This problem along with the use of sparse tensors to alleviate it are discussed by Zhao et al. (2023a).
@@
Generative networks were introduced to attack collaborative learning in (Hitaj et al., 2017; Wang et al., 2019a). However, contrary to optimization-based attacks and linear layer leakage, the attack does not aim to directly reconstruct exact training images but instead reconstructs representations of a targeted class which the attacker does not hold any images of. In the attack, an adversary (a client in collaborative or federated learning training) trains a generative network in conjunction with the benign collaborative model. For a single training round, the attack process goes as follows:
@@
In Wen et al. (2022), attackers attack a target image in an arbitrarily large batch by manipulating the model parameters to magnify the update of only the single image. Magnification is done in two steps: the magnification of the target class followed by a specific feature for a target image. In order to estimate a boundary for the features, the server uses several rounds of additional FL training for setup. This requires for the batch data to be the same across training rounds. Kariyappa et al. (2023) formulates the data reconstruction attack as a blind source separation problem. Given the observation that inputs to a linear layer can be directly reconstructed through the gradients, solving for the inputs is similar to solving for a unknown variables given a set of linear equations.
These attacks are all capable of breaking some levels of aggregation. (Pasquini et al., 2022; Lam et al., 2021) directly break aggregation and other data reconstruction attacks to work on top of them. (Wen et al., 2022; Kariyappa et al., 2023) both function on individual and aggregate updates. Each of these methods have their own strengths and weaknesses. Pasquini et al. (2022) requires sending different models to different clients but allows an attacker to single out a single update. Lam et al. (2021) requires multiple training rounds and additional knowledge in addition to the client update. However, it can essentially disaggregate aggregate updates over time. Kariyappa et al. (2023) can attack batch sizes up to 1024, but also requires malicious modification of the network parameters. Wen et al. (2022) targets a single image in an arbitary update size, but requires malicious modification of network parameters along with a few rounds of setup.